# NeuroGenPoisoning: Neuron-Guided Attacks on Retrieval-Augmented Generation of LLM via Genetic Optimization of External Knowledge

### v1

@@ 1 Introduction
Figure 1: An overview of NeuroGenPoisoning. This attack framework generates adversarial external knowledge in RAG guided by LLM internal neuron attribution and genetic optimization. Considering the internal-external knowledge conflict, it enables the generation of a large pool of effective poisoned knowledge at scale.
@@ 3.2 Poison-Responsive Neuron
This approach identifies neurons that are repeatedly sensitive to contextual information and uses them as a set of fixed targets to guide adversarial external knowledge generation. To further ground our approach theoretically, we provide a formal derivation in Appendix B, showing that maximizing the activation of Poison-Responsive Neurons directly increases the model’s predicted probability of the adversarial answer. This connection underscores the effectiveness of using neuron attribution as a principled fitness signal to optimize external knowledge.
@@ 4.1 Experimental Setup
To comprehensively assess the generalizability and effectiveness of our attack, we evaluate across three widely used open-domain QA datasets: SQuAD 2.0[30], TriviaQA[19], and WikiQA[46]. Each contains questions from diverse topics such as history, science, technology, sports, popular culture, and so on. Additional details are provided in Appendix A.
@@ 4.4.1 Effectiveness Across Models and Datasets
To evaluate the generalizability of our method, we measure POSR over multiple generations on three benchmark datasets: SQuAD 2.0, TriviaQA, and WikiQA, and across four open-source large language models: LLaMA-2-7b-chat-hf, Vicuna-7b-v1.5, Vicuna-13b-v1.5, and Gemma-7b. In our primary experiments, we set the number of Poison-Responsive Neurons $r$ = 10. This choice is based on the balance between computational efficiency and empirical performance. To assess whether our results are sensitive to this setting, we conducted a series of experiments with different $r$. We find that the final POSR remains stable in different settings. The details are shown in Appendix C.
@@
We also conducted a stratified analysis on the three datasets, grouping queries into the following knowledge domains: history & geography, literature, science & technology, and popular culture. For each domain, we measured POSR using the same global top Poison-Responsive Neurons set. We observe that POSR remains high (above 90%) in different knowledge domains and models, indicating that the attack is not limited to any specific domain or query structure. The details of the stratified analysis are shown in Appendix E.
@@ 6 Conclusion
Appendix A Datasets Descriptions
@@
Appendix B Theoretical Justification of Neuron-Guided Poisoning
@@
Appendix C Exploration of the Number of Poison-Responsive Neurons
@@
Appendix D Examples of Optimism External Knowledge and Output
@@
Appendix E Evaluation across Different Knowledge Domains
@@
Appendix F Details of Perplexity Evaluation
@@
Appendix G Broader Impacts

### v2

@@ 1 Introduction
Figure 1: Attack surfaces of RAG systems where malicious external content can override the model’s internal knowledge.
@@
Figure 2: The workflow of attack against RAG systems based on Poison-Responsive Neurons.
@@ 3.2 Poison-Responsive Neuron
This approach identifies neurons that are repeatedly sensitive to contextual information and uses them as a set of fixed targets to guide adversarial external knowledge generation. To further ground our approach theoretically, we provide a formal derivation in Appendix C, showing that maximizing the activation of Poison-Responsive Neurons directly increases the model’s predicted probability of the adversarial answer. This connection underscores the effectiveness of using neuron attribution as a principled fitness signal to optimize external knowledge.
@@ 4.1 Experimental Setup
To comprehensively assess the generalizability and effectiveness of our attack, we evaluate across three widely used open-domain QA datasets: SQuAD 2.0[30], TriviaQA[19], and WikiQA[46]. Each contains questions from diverse topics such as history, science, technology, sports, popular culture, and so on. Additional details are provided in Appendix B.
@@ 4.4.1 Effectiveness Across Models and Datasets
To evaluate the generalizability of our method, we measure POSR over multiple generations on three benchmark datasets: SQuAD 2.0, TriviaQA, and WikiQA, and across four open-source large language models: LLaMA-2-7b-chat-hf, Vicuna-7b-v1.5, Vicuna-13b-v1.5, and Gemma-7b. In our primary experiments, we set the number of Poison-Responsive Neurons $r$ = 10. This choice is based on the balance between computational efficiency and empirical performance. To assess whether our results are sensitive to this setting, we conducted a series of experiments with different $r$. We find that the final POSR remains stable in different settings. The details are shown in Appendix D.
@@
We also conducted a stratified analysis on the three datasets, grouping queries into the following knowledge domains: history & geography, literature, science & technology, and popular culture. For each domain, we measured POSR using the same global top Poison-Responsive Neurons set. We observe that POSR remains high (above 90%) in different knowledge domains and models, indicating that the attack is not limited to any specific domain or query structure. The details of the stratified analysis are shown in Appendix F.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported by University of Massachusetts Dartmouth MUST VI Research Program Funding. We thank all the reviewers and committee members for their valuable comments.
Appendix A Overview of the Framework
Our attack framework generates adversarial external knowledge in RAG guided by LLM internal neuron attribution and genetic optimization. Considering the internal-external knowledge conflict, it enables the generation of a large pool of effective poisoned knowledge at scale.
Figure 8: An overview of NeuroGenPoisoning.
Appendix B Datasets Descriptions
@@
Appendix C Theoretical Justification of Neuron-Guided Poisoning
@@
Appendix D Exploration of the Number of Poison-Responsive Neurons
@@
Appendix E Examples of Optimism External Knowledge and Output
@@
Appendix F Evaluation across Different Knowledge Domains
@@
Appendix G Details of Perplexity Evaluation
@@
Appendix H Broader Impacts